# Powering In-Database Dynamic Model Slicing for Structured Data Analytics

### v1

@@ 6.3. Ablation Study
In this subsection, we shall conduct an ablation study to evaluate the effectiveness of each component in LEADS.
@@
SQL-aware gating network. In this evaluation, we compare LEADS with two methods: w/o LEADS and LEADS w/o SQL-aware gating network. For the latter, we create a default SQL query vector by concatenating a set of default values for each attribute, denoted as $\textbf{q}_{d}=[\Delta_{1},\Delta_{2},\cdots,\Delta_{M}]$, indicating the absence of predicates in the SQL query for slicing a model. The comparison results are shown in Figure 8. There are two main observations. First, the w/o LEADS method achieves the lowest Workload-AUC because it simply uses the base model to handle the SQL queries. Second, the LEADS w/o SQL-aware gating network method results in a performance reduction across four base models compared to LEADS. For example, on the Credit dataset, this reduction can reach up to 0.02 in terms of Workload-AUC. It is because without the SQL-aware gating network, LEADS loses the ability for dynamic model customization based on SQL query vectors, leading to unsatisfactory results.
$\alpha$-entmax. In this investigation, we evaluate the effect of the $\alpha$-entmax function in LEADS. We compare LEADS to w/o LEADS and LEADS w/o $\alpha$-entmax, where the latter is a variant that substitutes the $\alpha$-entmax function with the softmax function. We use DNN as the base model and vary the number of experts from 2 to 256 to evaluate the performance w.r.t. Workload-AUC and FLOPs on Payment and Credit datasets.
@@
The results are presented in Figures 9-10. We can observe from Figure 9 that as the number of experts increases from 2 to 32, there is a notable improvement in AUC. This is expected because the model is able to generate more accurate predictions with additional experts. However, when the number of experts exceeds 32, there is a reduction in the Credit dataset for LEADS w/o $\alpha$-entmax, because the model becomes overly complex and results in overfitting and subsequently a decline in AUC performance. Figure 10 highlights the advantages of $\alpha$-entmax in terms of FLOPs saving, particularly when the number of experts increases. Specifically, the FLOPs of LEADS w/o $\alpha$-entmax show a linear increase with the growing number of experts. In contrast, LEADS with $\alpha$-entmax experiences an initial increase with a much gentler incline due to the fact that $\alpha$-entmax assigns small values to zero instead of retaining all experts when using softmax. In the slicing process, we then remove these unused experts with zero weights, effectively conserving computational resources.
@@
Figure 12. Analysis of the activated frequency for each expert w.r.t. the query workload. The x-axis denotes the expert ID, and the y-axis denotes the frequency (1.0 means that the expert is activated in every SQL query).
Regularization terms. For the balance term $L_{baln}$ and the sparsity term $L_{sprs}$, we compare the performance of LEADS with three variants: without the balance term (LEADS w/o $L_{baln}$), without the sparsity term (LEADS w/o $L_{sprs}$), and without both terms (LEADS w/o both). We use DNN as the base model and conduct experiments on the Payment and Credit datasets.
Figure 11 presents the comparison results w.r.t. Workload-AUC, and Figure 12 provides a breakdown analysis of activated frequency for each expert during the execution of the query workload, where a higher frequency indicates more extensive usage of an expert in the workload. There are three main findings from the results. First, removing the balance term significantly reduces the Workload-AUC of LEADS, as shown in Figure 11. While the sparse term is designed to counteract the effect of the balance term, using it solely will lead to a much smaller number of experts being utilized for each SQL query, resulting in lower prediction accuracy. This phenomenon can be validated in Figure 12, where only two experts are predominately selected. Second, when only adding the balance term, the performance is slightly lower than that of LEADS (see Figure 11), but the model utilizes almost all experts for every SQL query (see Figure 12), increasing computational costs. This is because the balance term encourages an even expert selection and drives the LEADS to utilize as many experts as possible. Third, when enabling both terms in LEADS, we can observe that the experts are utilized in a balanced manner while achieving the best performance, which demonstrates the effectiveness of our regularization terms.
(a) Response time for predicting 100k records on four datasets.
@@
Figure 13. Efficiency evaluation of INDICES in terms of response time. In the sub-figures, the ‘Left Bar’ denotes INDICES-decoupled (I-d), and the ‘Right Bar’ denotes INDICES (I).
@@ 6.4. System Efficiency
In this subsection, we evaluate the efficiency of INDICES in terms of its end-to-end response time and compare it with the INDICES-decoupled baseline (see Section 6.1.3). We also measure the breakdown response time for each system.
@@
Comparison with the baseline. We utilize a SQL query that selects 100k records for inference. We report the response time of INDICES and INDICES-decoupled on the four datasets, as presented in Figure 13(a). Compared to INDICES-decoupled, INDICES achieves speedup of 1.94x, 2.06x, 2.00x, and 1.82x on the Census, Credit, Diabetes, and Payment, respectively. There are three main reasons for such superior performance. First, INDICES reduces the costly data movement overheads between PostgreSQL and the inference system, with lower data retrieval time usage. Second, INDICES is further enhanced with the aforementioned optimizations: shared memory to reduce data copying overhead, and state caching to eliminate the cost of model loading during the entire inference UDF execution process. Last, within Frame’s Python execution, data from shared memory is directly read as a numpy.ndarray(Harris et al., 2020) , which facilitates quicker conversion to tensors for inference tasks and results in lower data preprocessing time in INDICES.
@@
Effects of the number of predicting records. We next evaluate the effect of the number of selected records in the SQL query in terms of the response time. We create SQL queries that select various numbers of records ranging from 40k to 640k from the Payment dataset. Figure 13(b) shows the response time of INDICES and INDICES-decoupled. We observe that INDICES consistently surpasses INDICES-decoupled across all record sizes, with performance improvement ranging from 1.47x to 1.93x. Moreover, the response time of INDICES increases more slowly than that of INDICES-decoupled. This is because the data movement overhead between the database and the inference system becomes more pronounced with more records.
@@
Evaluation of optimization techniques in INDICES. Further, we experiment to evaluate the benefits of the optimizations in Section 5.1. Specifically, we compare INDICES with (i) INDICES without memory sharing; (ii) INDICES without SPI; (iii) INDICES without state caching; and (iv) INDICES without any optimizations. Figure 14 presents the comparison results w.r.t. the response time to predict 100k records on the Payment dataset. The absence of shared memory results in significant data copying overhead between Rust and Python execution environments. Likewise, without SPI, the data retrieval time is high. Besides, if we do not enable state caching, it results in a substantial model loading overhead. With all the optimizations enabled, INDICES can greatly reduce the in-database inference response time.
@@ 7. Related Work
Mixture-of-Experts (Jacobs et al., 1991; Eigen et al., 2013; Yuksel et al., 2012) integrates the outputs of different experts in an input-driven manner. In the case of Sparse MoE, only a small subset of experts is chosen for each input, facilitating substantial model scaling without additional computation overhead. Sparse MoE has been used to build large language models (Shazeer et al., 2017; Fedus et al., 2022; Lepikhin et al., 2020) and applied to vision-related tasks(Riquelme et al., 2021; Xue et al., 2022). Our research delves into the potential of MoE in structured data analytics. We closely combine it with database data analytics, selecting experts based on the filter conditions in SQL queries.
In-Database Machine Learning involves running machine learning directly within the database. MADlib (Hellerstein et al., 2012) is an open-source library providing SQL-based ML functions in PostgreSQL. Google ML library(Goo, [n.d.]), and Microsoft’s SQL Server Machine Learning Services (mic, 2023) offer SQL APIs for ML functions on Oracle, bigquery, and Microsoft SQL Server, respectively. Recently,  (Xing et al., [n.d.]) proposed to integrate neural architectural search (NAS) model selection into PostgreSQL as an extension, which is orthogonal to our proposal. In summary, these works incorporate existing ML algorithms into the database for analysts but typically lack optimization for specific data analysis scenarios and seldom support deep learning algorithms.
Figure 14. Effects of INDICES’ optimization techniques.
@@ 8. Conclusions
In this paper, we propose a novel SQL-aware dynamic model slicing technique called LEADS. We enhance the general model trained on the entire database with the Mixture of Experts (MoE) technique and devise a SQL-aware gating network to effectively customize a sliced model given the propositional formula in the user’s SQL query. Further, we integrate LEADS into our end-to-end in-database inference system INDICES. We build the system on top of a full-fledged and open-source RDBMS, PostgreSQL, and introduce three optimization techniques to reduce the response time of inference queries. Extensive experiments conducted on four real-world datasets demonstrate that LEADS consistently outperforms four baseline models and INDICES significantly reduces the inference time compared to the conventional approach.

### v2

@@ 5.3. Ablation Study
In this part, we conduct the ablation study to answer the question RQ2, evaluating the effectiveness of each component in LEADS.
@@
SQL-aware gating network. In this evaluation, we remove the SQL-aware gating network to demonstrate the importance of dynamic expert selection. A default model is created using a special SQL query embedding, where a set of padding values for each attribute, denoted as $\textbf{q}_{d}=[\Delta_{1},\Delta_{2},\cdots,\Delta_{M}]$, indicating the absence of predicates in the SQL query. The comparison results are shown in Figure 8. There are two main observations. First, the w/o LEADS method achieves the lowest Workload-AUC because it simply uses the base model to handle all SQL queries. Second, the LEADS w/o SQL-aware gating network method results in a performance reduction compared to LEADS. For example, on the Credit dataset, this reduction can reach up to 0.02 in terms of Workload-AUC. It is because without the SQL-aware gating network, LEADS loses the ability for dynamic model customization based on SQL query vectors, leading to unsatisfactory results.
@@
$\alpha$-entmax. We evaluate the effect of the $\alpha$-entmax function by comparing LEADS to w/o LEADS and LEADS w/o $\alpha$-entmax (which uses the softmax function) Using DNN as the base model, we vary the number of experts from 2 to 256 and measure the performance w.r.t. Workload-AUC and FLOPs on Payment and Credit datasets. As shown in Figures 9, increasing the number of experts from 2 to 256 leads to notable AUC improvements, as more experts allow the model to make more accurate predictions. However, beyond 32 experts, LEADS w/o $\alpha$-entmax sees a decline in AUC on the Credit dataset due to easily overfitting, as the model becomes too complex. Figure 9 also highlights the advantages of $\alpha$-entmax in terms of FLOPs saving. Specifically, the FLOPs of LEADS w/o $\alpha$-entmax grow linearly with more experts, while LEADS with $\alpha$-entmax has a gentler increase, because $\alpha$-entmax assigns small values to zero, reducing the number of active experts compared to softmax. We remove these unused experts to conserve computational resources.
@@
Figure 11. Analysis of the activated frequency for each expert. The value denotes the frequency and 1.0 means the expert is activated in every SQL query.
(a) Response time for predicting 100k records on five datasets.
@@
(c) Response time comparison of SPJ and Non-SPJ queries.
(d) Example of SPJ (red box) and Non-SPJ queries (blue box).
Figure 12. Efficiency evaluation of In-database Inference Strategy (IIS) and Inference-Decouple Strategy (IDS).
Figure 13. Effects of optimizations on response time.
Regularization terms. We compare the performance of LEADS with three variants: without the balance term (LEADS w/o $L_{baln}$), without the sparsity term (LEADS w/o $L_{sprs}$), and without both terms (LEADS w/o both). Using DNN as the base model, we conduct experiments on the Payment and Credit datasets. Figure 10 presents the comparison results in w.r.t. Workload-AUC, and Figure 11 analyzes the frequency of expert usage during query execution. Three main findings emerge: First, removing the balance term significantly reduces the Workload-AUC of LEADS, Without the balance term, fewer experts are used for each query, leading to lower prediction accuracy. This is evident in Figure 11, where only two experts are predominately selected. Second, adding solely the balance term results in lower performance than LEADS but utilizes almost all experts for every query. The balance term encourages even expert selection, leading to higher computational costs. Lastly, enabling both terms simultaneously in LEADS balances expert usage while achieving the best performance.
@@ 5.4. System Efficiency
To answer question RQ3, we assess the system-level efficiency in end-to-end response time by comparing the IIS that is used in our system with IDS. We implement IIS as an extension that can be installed in PostgreSQL 14.
@@
Comparison with the baseline. Given an SQL query that selects 100k records for inference, We report the response time of IIS and IDS on the five datasets, as presented in Figure 12(a). Compared to IDS, IIS achieves a speedup of 1.94x, 2.06x, 2.00x, 1.82x, and 1.53x on the Census, Credit, Diabetes, Payment, and Avazu. There are two reasons for such superior performance. One, IIS reduces the costly data movement overheads between PostgreSQL and the inference system, with lower data retrieval time usage. Secondly, IIS is further enhanced with the optimizations: shared memory to reduce data copying overhead, and state caching to eliminate the cost of model loading during the inference UDF execution process.
@@
Effects of the number of predicting records. Next, we examine how the number of predictive records in the SQL query impacts response time. In the Payment dataset, this number ranges from 40k to 640k records. Figure 12(b) shows the response time for two strategies, IIS and IDS. We observe that IIS consistently surpasses IDS across various record numbers, with performance improvement ranging from 1.47x to 1.93x. Moreover, the response time of IIS increases more slowly than that of IDS, since the data movement overhead between the database and the inference system becomes more pronounced with the increase of records. Therefore, with more predictive records, IIS performs even more favorably.
@@
Evaluation of optimization techniques Further, we evaluate the benefits of the optimizations mentioned in Section 4.1. Specifically, we compare the in-database inference process with (i) w/o memory sharing; (ii) w/o SPI; (iii) w/o state caching; and (iv) w/o all optimizations. Figure 13 presents the comparison results w.r.t. the response time in predicting 100k records on the Payment dataset. The absence of shared memory leads to significant overhead due to data copying between different execution environments. Likewise, without SPI (PostgreSQL’s built-in data access API), data retrieval times are considerably longer. Moreover, there is considerable overhead from repeatedly loading the model without state caching. When all optimizations are enabled, the in-database inference extension achieves a 3x speed improvement compared to it without optimizations. Additionally, we investigate the trade-offs between response time and memory usage brought by memory sharing and state caching. In this analysis, we set a workload containing 300 inference queries and record memory usage during the process. The results presented in Figure 14 reveal that state caching and memory sharing enhance inference efficiency, with minimal impact on overall memory usage from these optimizations.
Figure 14. Optimizing UDFs: analyzing the trade-offs between response time and memory usage.
5.5. Complex Inference Scenarios
In this subsection, we address question RQ4 by examining LEADS’s effectiveness under data updates and schema changes, as well as the efficiency of our in-database inference extension when handling SPJ (Select-Project-Join) queries.
Effectiveness when data updates. We evaluate the effectiveness of LEADS for tuple insertion and deletion by executing the same SQL queries between data changes. Using the Credit and Diabetes datasets with a DNN as the base model, each dataset is equally divided into original and new tuples. The new tuples are inserted into the database in five stages, with the model evaluated after each insertion. As shown in Figure 15, the model enhanced by LEADS consistently outperforms the base model as the dynamic combination of multiple experts discussed in Section 5.2. In conclusion, when data updates follow the same distribution, the model enhanced by LEADS effectively generalizes to various subdatasets.
Effectiveness when the schema changes. As for schema changes, altering attributes changes the feature dimension of each tuple, leading to a dimension mismatch with the predictive model. To continue using the original model, we need to construct matched inputs. For insertion, required attributes can be fetched and fed to the model, while newly added attributes will not be considered in predictive modeling. For deletion, the values in the deleted attributes will be missing. We pad these missing values, but this will degrade performance. The impact of deleted attributes on model performance is shown in Figure 16. We notice that as more attributes are deleted, model performance drops rapidly. Therefore, retraining the model based on the new feature dimensions is advisable to ensure effective inference following schema changes.
Efficiency for SPJ queries. To evaluate in-database inference on SPJ queries, we vertically split the Diabetes and Credit datasets into two sub-tables (e.g., diabetes_left and diabetes_right). During inference, these sub-tables are joined to select tuples, as shown in the SQL example in Figure 12(d), and the efficiency is compared to the baseline approach IDS. We also assess performance using the original table without JOIN to measure the impact of query complexity. As expected, Figure 12(c) shows that JOIN operations introduce additional overhead, leading to higher response times compared to queries without JOIN. However, IDS outperforms the baselines due to reduced data movement, as data remains within the database. For JOIN queries, the speed advantage of in-database inference is less pronounced since query planning and execution times dominate over the benefit of eliminating network latency.
(a) Credit.
(b) Diabetes.
Figure 15. Effects of data insertion on Workload-AUC.
@@ 6. Related Work
Mixture-of-Experts (MoE) is initially proposed by (Jordan and Jacobs, 1994) to handle different samples using independent expert modules. (Shazeer et al., 2017) introduces the Sparse Gated MoE into language model training, developing large-scale LSTM-based MoE models, in which only one expert is chosen for each input data. With the rise of the Transformer as the dominant NLP architecture, researchers began to incorporate MoE layers by extending the Feed-Forward Networks (FFNs) in Transformers to build MoE language models. Despite these innovations, the Sparse Gated MoE (Shazeer et al., 2017) struggled with stability due to the expert route strategy where it brute-forcely selects the top-1 expert. To mitigate this, various studies (Lepikhin et al., 2020; Puigcerver et al., 2023; Fedus et al., 2022; Zhou et al., 2022) explored different learnable routing strategies for managing experts and input tokens. For example, Fast MoE (He et al., 2021) monitors the training status and dynamically adjusts the load for each expert. Expert Choice Routing (Zhou et al., 2022) lets experts choose tokens rather than selecting experts to prevent under-training. Soft MoE (Puigcerver et al., 2023) performs an implicit soft assignment by passing different weighted combinations of all input tokens to each expert. In addition to these works on MoE architectures and training strategies, recent years have witnessed the application of MoE in vision-related and multimodal predictive tasks (Riquelme et al., 2021). However, our research delves into the potential of MoE in structured data analytics. We closely combine it with database data analytics, dynamically selecting necessary experts corresponding to the filter conditions in SQL queries.
In-Database Machine Learning involves executing machine learning within the database. MADlib (Hellerstein et al., 2012) is an open-source library providing SQL-based ML functions in PostgreSQL. Google ML library(Goo, 2023), and Microsoft’s SQL Server Machine Learning Services (mic, 2023) offer SQL APIs for ML functions on Oracle, bigquery, and Microsoft SQL Server, respectively. However, they only support feeding the static model with tuples in a data table and do not adapt feeding it with meta-information such as query encoding to customize the model for inference. Adding this support requires significant changes to their infrastructure, which is either non-trivial or unfeasible since infrastructure code is not accessible. Therefore, none of them are directly comparable. Complementary to LEADS, we propose an effective in-database model selection technique, TRAILS, in  (Xing et al., 2024), and we will report their integration in the future.
(a) Credit.
(b) Diabetes.
Figure 16. Effects of schema deletion on Workload-AUC.
@@ 7. Conclusions
In this paper, we propose a novel SQL-aware dynamic model slicing technique called LEADS. We enhance the general model with the Mixture of Experts (MoE) technique and devise a SQL-aware gating network to dynamically customize a sliced model given the propositional formula in the user’s SQL query. We further encapsulate LEADS into an in-database inference extension for PostgreSQL. In the implementation, we incorporate three key optimizations to accelerate the in-database inference process. Extensive experiments on five real-world datasets show that LEADS consistently outperforms four baseline models, and the in-database inference extension significantly reduces inference time compared to traditional approaches. We have integrated LEADS into NeurDB, our ongoing implementation of an AI-powered autonomous data system.
Acknowledgement:
This research work is supported by Singapore Ministry of Education Academic Research Fund Tier 3 under MOE’s official grant number MOE2017-T3-1-007. Yuncheng Wu’s work is supported by National Key Research and Development Program of China (Grant No.2023YFB4503600).